# Vevo2: A Unified and Controllable Framework for Speech and Singing Voice Generation

### v2

@@ V Results and Analysis
Model
#Hz
Data
(#hours)
PT?
Expressive Speech Singing Voice
WER SIM
N-
CMOS
$S$-
CMOS
WER SIM
N-
CMOS
$S$-
CMOS
Ground Truth - - - 10.91 0.687 1.47 $_{\scriptscriptstyle\pm\text{0.24}}$ 0.87 $_{\scriptscriptstyle\pm\text{0.19}}$ 12.65 0.658 1.27 $_{\scriptscriptstyle\pm\text{0.14}}$ 0.55 $_{\scriptscriptstyle\pm\text{0.16}}$
F5-TTS [f5tts] - 101K 11.77 0.695 -1.05 $_{\scriptscriptstyle\pm\text{0.18}} -$0.06 $_{\scriptscriptstyle\pm\text{0.17}}$ 16.12 0.597 -1.89 $_{\scriptscriptstyle\pm\text{0.07}} -$1.29 $_{\scriptscriptstyle\pm\text{0.10}}$
MaskGCT [maskgct] 50 101K 13.42 0.736 -1.14 $_{\scriptscriptstyle\pm\text{0.21}}$ 0.07 $_{\scriptscriptstyle\pm\text{0.31}}$ 11.71 0.753 -1.74 $_{\scriptscriptstyle\pm\text{0.14}} -$0.97 $_{\scriptscriptstyle\pm\text{0.13}}$
CosyVoice 2 [cosyvoice] 25 167K 11.20 0.706 0.10 $_{\scriptscriptstyle\pm\text{0.29}} -$0.16 $_{\scriptscriptstyle\pm\text{0.13}}$ 16.18 0.659 -1.68 $_{\scriptscriptstyle\pm\text{0.15}} -$1.12 $_{\scriptscriptstyle\pm\text{0.16}}$
Vevo2-base
12.5
101K 15.52 0.677 -0.84 $_{\scriptscriptstyle\pm\text{0.15}} -$0.06 $_{\scriptscriptstyle\pm\text{0.27}}$ 19.39 0.658 -1.02 $_{\scriptscriptstyle\pm\text{0.16}} -$0.48 $_{\scriptscriptstyle\pm\text{0.18}}$
@@
Vevo2 12.5 101K, 7K 11.48 0.689 0.00 $_{\scriptscriptstyle\pm\text{0.00}}$ 0.00 $_{\scriptscriptstyle\pm\text{0.00}}$ 7.66 0.725 0.00 $_{\scriptscriptstyle\pm\text{0.00}}$ 0.00 $_{\scriptscriptstyle\pm\text{0.00}}$
    #Hz: Frame rate of the content-style tokenizer. Data: Pre-training data, where blue and red denote speech and singing data. The results on the regular speech domain are provided in the supplementary materials. The best and the second best result (excluding ground truth) are shown in bold and by underlined.
@@ V-A Benefits of Unified Modeling for Speech and Singing Voice
To investigate whether speech and singing voice generation can be mutually beneficial through unified modeling, we examine the impact of different pre-training data distributions on Vevo2. Specifically, we conduct experiments on zero-shot TTS tasks and evaluate the performance using reference prompts from different distributions. For comparison, we select models with varying content-style tokenizer frame rates: MaskGCT [maskgct] (50 Hz), CosyVoice 2 [cosyvoice2] (25 Hz), and F5-TTS [f5tts] (a single-stage model without content-style tokenizer). The results are presented in Table I, which demonstrates the following findings:
@@ V-B Controllability over Text and Prosody
Model English Chinese
WER SIM FPC
N-CMOS
Melody-MOS
WER SIM FPC
N-CMOS
Melody-MOS
Ground Truth 17.89 0.640 0.815 - $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}}$ 7.40 0.675 0.782 - $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}}$
StyleSinger [stylesinger] - - - - $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}}$ 14.04 0.555 0.670 -0.37 $_{\scriptscriptstyle\pm\text{0.16}}$ 2.33 $_{\scriptscriptstyle\pm\text{0.36}}$
TCSinger [tcsinger] 40.83 0.605 0.724 -0.09 $_{\scriptscriptstyle\pm\text{0.12}}$ 2.55 $_{\scriptscriptstyle\pm\text{0.27}}$ 14.26 0.651 0.707 -0.28 $_{\scriptscriptstyle\pm\text{0.17}}$ 2.58 $_{\scriptscriptstyle\pm\text{0.19}}$
Vevo2-base 60.83 0.586 0.719 - $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}}$ 23.60 0.653 0.674 - $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}}$
Vevo2 24.77 0.643 0.752 0.00 $_{\scriptscriptstyle\pm\text{0.00}}$ 2.29 $_{\scriptscriptstyle\pm\text{0.26}}$ 9.83 0.669 0.710 0.00 $_{\scriptscriptstyle\pm\text{0.00}}$ 2.44 $_{\scriptscriptstyle\pm\text{0.18}}$
@@
Model Expressive Speech Singing Voice
@@
SSR-Speech [ssrspeech] 29.56 0.588 0.721 -1.36 $_{\scriptscriptstyle\pm\text{0.09}} -$0.89 $_{\scriptscriptstyle\pm\text{0.14}}$ 51.92 0.710 0.804 -1.43 $_{\scriptscriptstyle\pm\text{0.17}} -$1.10 $_{\scriptscriptstyle\pm\text{0.08}}$
F5-TTS [f5tts] 21.35 0.733 0.730 -0.19 $_{\scriptscriptstyle\pm\text{0.12}} -$0.21 $_{\scriptscriptstyle\pm\text{0.16}}$ 29.78 0.784 0.821 -1.15 $_{\scriptscriptstyle\pm\text{0.23}} -$0.97 $_{\scriptscriptstyle\pm\text{0.19}}$
Vevo2-base 23.54 0.795 0.782 - $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}}$ 29.69 0.841 0.872 - $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}}$
Vevo2 16.83 0.799 0.792 0.00 $_{\scriptscriptstyle\pm\text{0.00}}$ 0.00 $_{\scriptscriptstyle\pm\text{0.00}}$ 17.98 0.848 0.877 0.00 $_{\scriptscriptstyle\pm\text{0.00}}$ 0.00 $_{\scriptscriptstyle\pm\text{0.00}}$
@@
Zero-Shot SVS Given a target text, a music notation like MIDI, and a reference singer’s waveform, a zero-shot SVS model is required to generate singing voice that matches the content of the target text and music notation while preserving the timbre and style of the reference singer [stylesinger, tcsinger, spsinger]. For Vevo2, we render MIDI into piano waveform and use it as the prosodic source, which can also be considered as instrument-to-singing. The results are shown in Table II, where we include StyleSinger [stylesinger] and TCSinger [tcsinger] as two baselines.
@@
Zero-Shot Editing In addition to SVS, we adopt editing tasks to evaluate Vevo2’s text and prosody controllability. The used evaluation sets are constructed with DeepSeek-V3 [deepseek-v3]. Specifically, we utilized the same expressive speech and singing voice evaluation data from our zero-shot TTS tasks. For each sample’s text, we generated edited versions using DeepSeek-V3 with the following prompt: "Please edit the following text by adding, deleting, or modifying content as needed. Note: (1) Ignore any existing punctuation errors in the original text and ensure proper punctuation in the edited version. (2) The edited version must differ from the original text. (3) Return only the final edited text without any additional explanations."
Vevo2’s advantage in editing tasks stems from its use of prosody tokens as explicit control, enabling it to preserve the prosody (melody) of the raw voice during editing. For instance, in singing lyric editing tasks [editsinger, unisinger], Vevo2 can modify only the lyrics while maintaining the original singing voice melody. The results are shown in Table II, where we include two speech editing baselines, SSR-Speech [ssrspeech] and F5-TTS [f5tts]. It can be observed that compared to the baselines, the post-trained Vevo2 not only achieves better intelligibility (WER), naturalness (N-CMOS), and speaker similarity (SIM), but also demonstrates superior prosody preservation of the original voice (PS-CMOS), particularly in the singing voice domain (MS-CMOS). We will provide more audio samples at our website to further demonstrate Vevo2’s advantages in singing lyric editing tasks.
Model English Chinese
WER SIM
N-
CMOS
Style-
CMOS
Melody-
MOS
WER SIM
N-
CMOS
Style-
CMOS
Melody-
MOS
Ground Truth 16.48 - - $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}}$ 12.82 - - $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}}$
FACodec [naturalspeech3] 32.81 0.434 - $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}}$ 36.97 0.473 - $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}}$
Vevo-FM [vevo] 24.05 0.567 -0.64 $_{\scriptscriptstyle\pm\text{0.07}} -$0.45 $_{\scriptscriptstyle\pm\text{0.16}}$ 0.71 $_{\scriptscriptstyle\pm\text{0.21}}$ 22.85 0.610 -0.55 $_{\scriptscriptstyle\pm\text{0.19}} -$0.44 $_{\scriptscriptstyle\pm\text{0.08}}$ 0.88 $_{\scriptscriptstyle\pm\text{0.27}}$
CosyVoice2-FM [cosyvoice2] 23.59 0.553 -0.64 $_{\scriptscriptstyle\pm\text{0.11}} -$0.63 $_{\scriptscriptstyle\pm\text{0.18}}$ 2.25 $_{\scriptscriptstyle\pm\text{0.15}}$ 20.14 0.589 -0.63 $_{\scriptscriptstyle\pm\text{0.15}} -$0.54 $_{\scriptscriptstyle\pm\text{0.07}}$ 1.65 $_{\scriptscriptstyle\pm\text{0.23}}$
NeuCoSVC 2 [neucosvc] 30.61 0.481 - $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}}$ 32.26 0.519 - $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}}$
SeedVC (SVC) [seedvc] 22.86 0.508 -0.05 $_{\scriptscriptstyle\pm\text{0.13}} -$0.24 $_{\scriptscriptstyle\pm\text{0.15}}$ 2.89 $_{\scriptscriptstyle\pm\text{0.09}}$ 15.65 0.550 -0.33 $_{\scriptscriptstyle\pm\text{0.21}} -$0.54 $_{\scriptscriptstyle\pm\text{0.14}}$ 2.93 $_{\scriptscriptstyle\pm\text{0.14}}$
Vevo2-FM 29.82 0.587 -0.12 $_{\scriptscriptstyle\pm\text{0.16}} -$0.11 $_{\scriptscriptstyle\pm\text{0.12}}$ 2.91 $_{\scriptscriptstyle\pm\text{0.12}}$ 22.54 0.611 -0.03 $_{\scriptscriptstyle\pm\text{0.30}} -$0.32 $_{\scriptscriptstyle\pm\text{0.17}}$ 2.90 $_{\scriptscriptstyle\pm\text{0.08}}$
Vevo2 11.64 0.601 0.00 $_{\scriptscriptstyle\pm\text{0.00}}$ 0.00 $_{\scriptscriptstyle\pm\text{0.00}}$ 2.24 $_{\scriptscriptstyle\pm\text{0.18}}$ 14.53 0.623 0.00 $_{\scriptscriptstyle\pm\text{0.00}}$ 0.00 $_{\scriptscriptstyle\pm\text{0.00}}$ 2.38 $_{\scriptscriptstyle\pm\text{0.25}}$
@@
Model Accent Conversion Emotion Conversion
@@
Vevo [vevo] 30.37 0.702 0.526 3.49 13.31 0.697 0.406 3.95
Vevo2-base 10.08 0.760 0.537 3.72 11.36 0.700 0.400 3.82
Vevo2 7.73 0.768 0.529 3.76 7.89 0.714 0.402 3.89
@@
Task DDUR ($\downarrow$)
DC ($\uparrow$)

### v3

@@ V Results and Analysis
Model #Hz Data (#hours) #Param PT? Expressive Speech Singing Voice
WER SIM N- CMOS SS- CMOS WER SIM N- CMOS SS- CMOS
Ground Truth - - - - 10.91 0.687 1.47 $_{\scriptscriptstyle\pm\text{0.24}}$ 0.87 $_{\scriptscriptstyle\pm\text{0.19}}$ 12.65 0.658 1.27 $_{\scriptscriptstyle\pm\text{0.14}}$ 0.55 $_{\scriptscriptstyle\pm\text{0.16}}$
F5-TTS [11] - 101K 336M 11.77 0.695 -1.05 $_{\scriptscriptstyle\pm\text{0.18}} -$0.06 $_{\scriptscriptstyle\pm\text{0.17}}$ 16.12 0.597 -1.89 $_{\scriptscriptstyle\pm\text{0.07}} -$1.29 $_{\scriptscriptstyle\pm\text{0.10}}$
MaskGCT [71] 50 101K 1048M 13.42 0.736 -1.14 $_{\scriptscriptstyle\pm\text{0.21}}$ 0.07 $_{\scriptscriptstyle\pm\text{0.31}}$ 11.71 0.753 -1.74 $_{\scriptscriptstyle\pm\text{0.14}} -$0.97 $_{\scriptscriptstyle\pm\text{0.13}}$
CosyVoice 2 [21] 25 167K 581M 11.20 0.706 0.10 $_{\scriptscriptstyle\pm\text{0.29}} -$0.16 $_{\scriptscriptstyle\pm\text{0.13}}$ 16.18 0.659 -1.68 $_{\scriptscriptstyle\pm\text{0.15}} -$1.12 $_{\scriptscriptstyle\pm\text{0.16}}$
Vevo2-base 12.5 101K 872M 15.52 0.677 -0.84 $_{\scriptscriptstyle\pm\text{0.15}} -$0.06 $_{\scriptscriptstyle\pm\text{0.27}}$ 19.39 0.658 -1.02 $_{\scriptscriptstyle\pm\text{0.16}} -$0.48 $_{\scriptscriptstyle\pm\text{0.18}}$
@@
Vevo2 12.5 101K, 7K 872M 11.48 0.689 0.00 $_{\scriptscriptstyle\pm\text{0.00}}$ 0.00 $_{\scriptscriptstyle\pm\text{0.00}}$ 7.66 0.725 0.00 $_{\scriptscriptstyle\pm\text{0.00}}$ 0.00 $_{\scriptscriptstyle\pm\text{0.00}}$
    #Hz: Frame rate of the content-style tokenizer. Data: Pre-training data, where blue and red denote speech and singing data. #Param: The number of learnable parameters. The results on the regular speech domain are provided in the supplementary materials. The best and the second best result (excluding ground truth) are shown in bold and by underlined.
@@ V-A Benefits of Unified Modeling for Speech and Singing Voice
To investigate whether speech and singing voice generation can be mutually beneficial through unified modeling, we examine the impact of different pre-training data distributions on Vevo2. Specifically, we conduct experiments on zero-shot TTS tasks and evaluate the performance using reference prompts from different distributions. For comparison, we select models with varying content-style tokenizer frame rates: MaskGCT [71] (50 Hz), CosyVoice 2 [22] (25 Hz), and F5-TTS [11] (a single-stage model without content-style tokenizer). The results are presented in Table I, which demonstrates the following findings:
@@ V-B Controllability over Text and Prosody
Model #Param English Chinese
WER SIM FPC N- CMOS Melody- MOS WER SIM FPC N- CMOS Melody- MOS
Ground Truth - 17.89 0.640 0.815 - $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}}$ 7.40 0.675 0.782 - $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}}$
StyleSinger [97] 42M - - - - $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}}$ 14.04 0.555 0.670 -0.37 $_{\scriptscriptstyle\pm\text{0.16}}$ 2.33 $_{\scriptscriptstyle\pm\text{0.36}}$
TCSinger [98] 330M 40.83 0.605 0.724 -0.09 $_{\scriptscriptstyle\pm\text{0.12}}$ 2.55 $_{\scriptscriptstyle\pm\text{0.27}}$ 14.26 0.651 0.707 -0.28 $_{\scriptscriptstyle\pm\text{0.17}}$ 2.58 $_{\scriptscriptstyle\pm\text{0.19}}$
Vevo2-base 872M 60.83 0.586 0.719 - $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}}$ 23.60 0.653 0.674 - $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}}$
Vevo2 872M 24.77 0.643 0.752 0.00 $_{\scriptscriptstyle\pm\text{0.00}}$ 2.29 $_{\scriptscriptstyle\pm\text{0.26}}$ 9.83 0.669 0.710 0.00 $_{\scriptscriptstyle\pm\text{0.00}}$ 2.44 $_{\scriptscriptstyle\pm\text{0.18}}$
@@
Model #Param Expressive Speech Singing Voice
@@
SSR-Speech [70] 830M 29.56 0.588 0.721 -1.36 $_{\scriptscriptstyle\pm\text{0.09}} -$0.89 $_{\scriptscriptstyle\pm\text{0.14}}$ 51.92 0.710 0.804 -1.43 $_{\scriptscriptstyle\pm\text{0.17}} -$1.10 $_{\scriptscriptstyle\pm\text{0.08}}$
F5-TTS [11] 336M 21.35 0.733 0.730 -0.19 $_{\scriptscriptstyle\pm\text{0.12}} -$0.21 $_{\scriptscriptstyle\pm\text{0.16}}$ 29.78 0.784 0.821 -1.15 $_{\scriptscriptstyle\pm\text{0.23}} -$0.97 $_{\scriptscriptstyle\pm\text{0.19}}$
Vevo2-base 872M 23.54 0.795 0.782 - $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}}$ 29.69 0.841 0.872 - $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}}$
Vevo2 872M 16.83 0.799 0.792 0.00 $_{\scriptscriptstyle\pm\text{0.00}}$ 0.00 $_{\scriptscriptstyle\pm\text{0.00}}$ 17.98 0.848 0.877 0.00 $_{\scriptscriptstyle\pm\text{0.00}}$ 0.00 $_{\scriptscriptstyle\pm\text{0.00}}$
@@
Model #Param English Chinese
WER SIM N- CMOS Style- CMOS Melody- MOS WER SIM N- CMOS Style- CMOS Melody- MOS
Ground Truth - 16.48 - - $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}}$ 12.82 - - $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}}$
FACodec [36] 139M 32.81 0.434 - $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}}$ 36.97 0.473 - $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}}$
Vevo-FM [95] 334M 24.05 0.567 -0.64 $_{\scriptscriptstyle\pm\text{0.07}} -$0.45 $_{\scriptscriptstyle\pm\text{0.16}}$ 0.71 $_{\scriptscriptstyle\pm\text{0.21}}$ 22.85 0.610 -0.55 $_{\scriptscriptstyle\pm\text{0.19}} -$0.44 $_{\scriptscriptstyle\pm\text{0.08}}$ 0.88 $_{\scriptscriptstyle\pm\text{0.27}}$
CosyVoice2-FM [22] 75M 23.59 0.553 -0.64 $_{\scriptscriptstyle\pm\text{0.11}} -$0.63 $_{\scriptscriptstyle\pm\text{0.18}}$ 2.25 $_{\scriptscriptstyle\pm\text{0.15}}$ 20.14 0.589 -0.63 $_{\scriptscriptstyle\pm\text{0.15}} -$0.54 $_{\scriptscriptstyle\pm\text{0.07}}$ 1.65 $_{\scriptscriptstyle\pm\text{0.23}}$
NeuCoSVC 2 [59] 14M 30.61 0.481 - $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}}$ 32.26 0.519 - $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}} -$ $_{\scriptscriptstyle\pm\text{0.00}}$
SeedVC (SVC) [46] 200M 22.86 0.508 -0.05 $_{\scriptscriptstyle\pm\text{0.13}} -$0.24 $_{\scriptscriptstyle\pm\text{0.15}}$ 2.89 $_{\scriptscriptstyle\pm\text{0.09}}$ 15.65 0.550 -0.33 $_{\scriptscriptstyle\pm\text{0.21}} -$0.54 $_{\scriptscriptstyle\pm\text{0.14}}$ 2.93 $_{\scriptscriptstyle\pm\text{0.14}}$
Vevo2-FM 363M 29.82 0.587 -0.12 $_{\scriptscriptstyle\pm\text{0.16}} -$0.11 $_{\scriptscriptstyle\pm\text{0.12}}$ 2.91 $_{\scriptscriptstyle\pm\text{0.12}}$ 22.54 0.611 -0.03 $_{\scriptscriptstyle\pm\text{0.30}} -$0.32 $_{\scriptscriptstyle\pm\text{0.17}}$ 2.90 $_{\scriptscriptstyle\pm\text{0.08}}$
Vevo2 872M 11.64 0.601 0.00 $_{\scriptscriptstyle\pm\text{0.00}}$ 0.00 $_{\scriptscriptstyle\pm\text{0.00}}$ 2.24 $_{\scriptscriptstyle\pm\text{0.18}}$ 14.53 0.623 0.00 $_{\scriptscriptstyle\pm\text{0.00}}$ 0.00 $_{\scriptscriptstyle\pm\text{0.00}}$ 2.38 $_{\scriptscriptstyle\pm\text{0.25}}$
@@
Zero-Shot SVS Given a target text, a music notation like MIDI, and a reference singer’s waveform, a zero-shot SVS model is required to generate singing voice that matches the content of the target text and music notation while preserving the timbre and style of the reference singer [97, 98, 101]. For Vevo2, we render MIDI into piano waveform and use it as the prosodic source, which can also be considered as instrument-to-singing. The results are shown in Table II, where we include StyleSinger [97] and TCSinger [98] as two baselines.
@@
Zero-Shot Editing In addition to SVS, we adopt editing tasks to evaluate Vevo2’s text and prosody controllability. The used evaluation sets are constructed with DeepSeek-V3 [18]. Specifically, we utilized the same expressive speech and singing voice evaluation data from our zero-shot TTS tasks. For each sample’s text, we generated edited versions using DeepSeek-V3 with the following prompt: "Please edit the following text by adding, deleting, or modifying content as needed. Note: (1) Ignore any existing punctuation errors in the original text and ensure proper punctuation in the edited version. (2) The edited version must differ from the original text. (3) Return only the final edited text without any additional explanations."
Vevo2’s advantage in editing tasks stems from its use of prosody tokens as explicit control, enabling it to preserve the prosody (melody) of the raw voice during editing. For instance, in singing lyric editing tasks [90, 31], Vevo2 can modify only the lyrics while maintaining the original singing voice melody. The results are shown in Table II, where we include two speech editing baselines, SSR-Speech [70] and F5-TTS [11]. It can be observed that compared to the baselines, the post-trained Vevo2 not only achieves better intelligibility (WER), naturalness (N-CMOS), and speaker similarity (SIM), but also demonstrates superior prosody preservation of the original voice (PS-CMOS), particularly in the singing voice domain (MS-CMOS). We will provide more audio samples at our website to further demonstrate Vevo2’s advantages in singing lyric editing tasks.
@@
Model #Param Accent Conversion Emotion Conversion
@@
Vevo [95] 797M 30.37 0.702 0.526 3.49 13.31 0.697 0.406 3.95
Vevo2-base 872M 10.08 0.760 0.537 3.72 11.36 0.700 0.400 3.82
Vevo2 872M 7.73 0.768 0.529 3.76 7.89 0.714 0.402 3.89
@@
Task DDUR ($\downarrow$) DC ($\uparrow$)